### v1

# Exoplanet Detection : A Detailed Analysis

Abstract
The exoplanet detection is the most exciting and challenging field of astronomy. The discovery of many exoplanets has revolutionized our understanding of the formation and evolution of planetary systems and has showed new ways to search for extra terrestrial life. In recent years, some primary methods of exoplanet detection like transit, radial velocity, gravitational microlensing, direct imaging and astrometry have played a important role for the discovery of exoplanets. In this paper we explored detection methodologies with all the implications and analytics of comparison between them. Here we also discussed on different machine learning algorithms for exoplanet detection and visualization. Finally, concluded with the significant discoveries made by some missions and their implications on our understanding for the properties, environmental conditions and importance of exoplanets in the universe.
keywords:
Exoplanet Detection, Transit, Missions, Exoplanets, Gaia
1 Introduction
Since the discovery of exoplanets began, they have continually surprised us with the hidden secrets of the cosmos. Here on earth, wealth is often associated with the possession of precious rocks and metals. But what about worlds made of these rocks and metals? For instance, 55 Cancri e is an exoplanet made of diamonds and orbits a sun-like star. It was discovered by McArthur McArthur et al. (2004) using the radial velocity method by observing the spectrum of one of the parent stars in a binary system located 40 light-years from the planet Earth. Exoplanets are poised to become a significant part of human civilization in various aspects, such as energy sources, colonization of other planets, transportation, space exploration, employment, and many others. Since the first exoplanet was discovered, astronomers have been working tirelessly to discover other extrasolar planets. Numerous missions have been conducted by several space agencies to study exoplanets, and the results have been truly fulfilling. Till date, approximately 5500+ exoplanets have been discovered Akeson et al. (2013), with many of them undergoing detailed studies.
Figure 1: Exoplanet detection methods, shown on a logarithmic mass scale, show current and upcoming measuring possibilities. Open arrows show previous evaluations, whereas solid arrows represent discoveries. Data collected from the NASA Exoplanet Archive (2018) Perryman (2018).
Exoplanets come in various types such as earth-size, neptune-size, super-earth, jupiter-size, sub-earth, super-jupiters, brown dwarfs, and stellar companions. There are also various techniques for detecting exoplanets Wei (2018), including radial velocity, transit method, astrometry, direct imaging, and gravitational microlensing. Each method offers distinct perspectives, strengths, limitations, and other details, often represented in tabular and graphical forms. By analyzing the light that passes through the atmosphere of a planet, scientists can understand its composition. Some exoplanets have been found to have atmospheres containing elements such as carbon dioxide, methane, and even water vapor Perryman (2018). If we compare a piece of wood and a piece of diamond to see which one is more expensive, the correct answer is wood. This is because, so far, earth is the only place in the universe where wood is produced, while diamonds are present on different exoplanets like 55 Cancri e. Gliese 667Cc is a super-earth exoplanet that lies 22 light-years away from earth. It’s a potentially habitable super-earth with a mass of 3.8 times that of earth. Humans are still in the race of hunting for those habitable exoplanets, on which we can grow more plants and expand human colonization beyond Earth.
51 Pegasi b, the first exoplanet, orbiting around a star similar to sun, was found using the radial velocity approach by Mayor & Queloz. Several space agencies are running missions specifically for exoplanet detection using different detection methods Shafi (2021). These missions include hubble, spitzer, corot, kepler/K2, gaia, TESS, CHEOPs, webb, and many more, which continuously monitor the universe, and astronomers keep detecting exoplanets with their help. In the future, there will be missions for specific exoplanets to understand their properties in detail and their environmental conditions. However, as of now, it’s hard to travel to even nearby exoplanets within a lifetime. Further more, introduced space missions briefly, including details about data formats and other details. Astronomers use various machine learning Kumari et al. (2023) and data cleaning algorithms to work on the data that comes from these missions. Such algorithms include the BLS algorithm, light curve analyzer, CNN, and other algorithms. all exoplanets are formed by normal matter but some scientists assumes that some exoplanets could me formed through dark matter Bai et al. (2023). Introduction about exoplanets with there basic details are going to be dicsussed in the next section.
2 The exoplanets
Extrasolar planets, also referred to as exoplanets, are planets that circle stars outside of our solar system. From the first exoplanet to be discovered in the early 1990s to the current booming field with thousands of confirmed planets, the growth of exoplanet investigation has been incredible wikipedia (2022). Section 2 discusses the crucial roles that various detection methods have had in this process. Technological developments like adaptive optics, space-based observatories like kepler and TESS, and high-precision spectrographs have made these approaches far more useful Woodrum et al. (2023). The discovery of smaller and farther-off planets has been made feasible by these developments. Exoplanet research is essential because it can increase our knowledge of planetary systems and the conditions necessary for habitability and the possibility that life exists elsewhere in the cosmos.
2.1 Stars That Host Planets
On average, each star Scalo et al. (2007) is believed to host at least one planet. It’s estimated that around 20% of sun like stars have an "earth-sized" planet in their habitable zones. The majority of known exoplanets orbit stars similar to our Sun, specifically those in the main-sequence spectral categories F, G, or K Lee et al. (2012) Perryman (2018) Meunier et al. (2023). Planets detectable by the radial-velocity method are less common around lower mass stars, such as the red dwarfs in the M spectral group Ananyeva et al. (2023). Despite this, the transit method, which excels at detecting smaller planets, has enabled the kepler telescope to identify dozens of planets orbiting near red dwarfs.
Figure 2: Diagram showing the habitable zones surrounding certain stars. The green area indicates where water on a planet’s surface can stay liquid at the ideal distance from the star. The blue area is too cool, and the red area is too heated.
2.2 Detection methods
Various detection methods Rouan & Lagrange (2023) exist for exoplanets, and astronomers continuously work toward developing new ways of detection. Some effective methods include radial velocity, which led to the discovery of the 1st exoplanet in the late ’90s; the transit method, which analyzes light curves and periodograms Bozza et al. (2016); direct imaging, capturing exoplanets through images; gravitational microlensing; and astrometry Perryman (2018). Table 1 shows the total number of exoplanets discovered by various methods of detection till beginning of 2024 NASA (2024). data taken from NASA exoplanet archive Akeson et al. (2013).
Table 1: Exoplanets discovered by various methods till date
Discovered Exoplanets
Method of discovery Number of planets
Astrometry 3
Disk Kinematics 1
Radial Velocity 1071
Eclipse timing variations 17
Orbital brightness modulations 9
Transit 4146
Transit timing variations 28
Pulsar timing variations 7
Microlensing 204
Imaging 69
Pulsation timing variations 2
2.3 Types of exoplanets
As of the conclusion of 2023, the existence of over 5500 planets had been confirmed. In an attempt to better understand these celestial bodies, scientists have undertaken numerous initiatives to categorize them based on various attributes, including their size, mass, and composition Perryman (2018).Table 2 shows the classification of exoplanets based on mass, size and chemical compositions of known gases such as $ZnS$, $H_{2}O$, $CO_{2}$ $CH_{4}$ Forestano et al. (2023).
Figure 3: planet masses are categorized based on solar system objects, as suggested by Stevens & Gaudi (2013). The solar system’s planet masses are displayed on the right, with the size of the circles corresponding to the planets’ radii Perryman (2018) Mathur et al. (2021).
Table 2: Types of exoplanets based on various factors
Factor Planet type Parameter
Size Earth size ($1.25-2R_{\oplus}$)
Super Earth size ($2-6R_{\oplus}$)
Jupiter size ($6-15R_{\oplus}$)
Mass Earth ($0.1M_{\oplus}-2M_{\oplus}$)
Sub Earth ($10^{-8}M_{\oplus}-0.1M_{\oplus}$)
Super Earth ($2M_{\oplus}-10M_{\oplus}$)
Neptunes ($10M_{\oplus}-100M_{\oplus}$)
Jupiters ($100M_{\oplus}-10^{3}M_{\oplus}$)
Super Jupiters ($10^{3}M_{\oplus}-13M_{J}$)
Brown Dwarfs ($13M_{J}-0.07M_{\odot}$)
Stellar Companions ($0.07M_{\odot}-1M_{\odot}$)
Composition Hot super Earths ($187-1.12$)
(stellar flux) Warm Super Earths ($1.12-0.30$)
Cold super Earth ($0.30-0.0030$)
Hot Rocky ($182-1.0$)
Warm Rocky ($1.0-0.28$)
Cold Rocky ($0.28-0.0035$)
Hot Sub Neptunes ($188-1.15$)
Warm Sub Neptunes ($1.15-0.32$)
Cold Sub Neptunes ($0.32-0.0.0030$)
Hot Sub Jovians ($220-1.65$)
Warm Sub Jovians ($1.65-0.45$)
Cold Sub Jovians ($0.45-0.0030$)
Hot Jovians ($220-1.65$)
Warm Jovians ($1.65-0.40$)
Cold Jovians ($0.40-0.0025$)
2.4 Atmospheric conditions
Finding transiting exoplanets is a significant breakthrough that provides previously unattainable knowledge about the composition and physical properties of these far-off worlds Yakovlev et al. (2022). Transit and eclipse spectroscopy allows the investigation of atmospheric dynamics Walker et al. (2018) and heat transport processes, especially for hot Jupiters, while the resulting densities, derived from detailed measurements of masses and radii, offer important hints regarding internal structures and compositions. These discoveries have been made possible in large part by drawing on knowledge amassed over the previous 50 years from research on the planets and satellites of the solar system Jakka (2023).
Future developments in photometric and spectroscopic observations, such as those brought about by the 2020 launch of the James Webb Space Telescope (JWST), are very promising. Advanced diagnostics may reveal the precise atmospheric compositions of exoplanets that are located in the "habitable zone Jagtap et al. (2021), " which is where liquid water may be present. These compositions may include indications of either primitive or advanced life forms. These developments advance not only the hunt for extraterrestrial life but also help answer the important question of how common or uncommon life is in the universe Seager & Deming (2010). The dynamic interplay between theoretical models and observational methods is set to propel future research and transform our knowledge of exoplanetary systems.
3 Exoplanet detection methods
As discussed previously there are various detection methods which played an important role to present this unique field of study. all methods works on various different properties such as light, noise Smith et al. (2008), EMwaves etc. some of them are explained here.
3.1 Radial Velocity Method
The radial velocity technique for identifying exoplanets involves tracking the reflexive movement of a star, which is triggered by the gravitational force of a planet in orbit. This method assesses three visible characteristics of the star that are altered by the planet’s existence:
Radial Velocity: When a planet orbits a star, both celestial bodies revolve around a shared center of mass, termed the barycenter. The star’s motion around this point results in a periodic fluctuation in its radial velocity Hatzes (2016), which signifies its speed moving towards or away from an observer located on Earth. It is discovered by spectroscopic studies that this radial velocity shift exists. The spectral lines of a star change towards the blue end of the spectrum as it travels towards the observer, and to the red end when it goes away (blueshift and redshift, respectively) Bozza et al. (2016) Chen (2023).
Angular (Astrometric) Position in the Sky: The star moves in a modest circular or elliptical path around the shared center of mass due to the gravitational attraction of an orbiting planet Pham et al. (2024) Bozza et al. (2016). The star’s angular location in the sky changes periodically, indicating its motion. Astrometric measures entail monitoring the star’s apparent location with accuracy over time. Measurements of radial velocity have been used more frequently than this approach, which is more difficult Perryman (2018) Chen (2023).
Time of Arrival of Periodic Reference Signal: Variations in the arrival times of periodic signals, such pulsar pulses or eclipse timings in binary star systems, can be caused by the gravitational interaction between a star and its circling planet Bozza et al. (2016). Timing variations in these signals of reference may indicate the existence of an unseen planet and offer additional evidence of its impact Fischer et al. (2015).
Radial velocity Dumusque et al. (2017) measurements revealed the first credible discovery of an exoplanet in 1995 Perryman (2018). Since then, a number of exoplanets have been found using this method because to technological and observational breakthroughs. Via radial velocity metrics, 1075 exoplanets had been found as of 02/05/2024; some of these planets have been observed to transit their host stars, allowing the star’s light to periodically fade as the planet passes in front of it.
3.2 Transit method
Astronomy achieved a major milestone in 1999 with the ground-breaking observation of the first transit of an extrasolar planet Perryman (2018) Chen (2023). The ’hot Jupiter’ system HD 209458 was observed photometrically Perryman (2018), which allowed for this observation. A gas giant planet known as a "hot Jupiter" has a close orbit around its host star, which increases the planet’s temperature somewhat. other star systems found by radial velocity surveys came into focus after this finding Fischer et al. (2015). After a planet passes in front of its host star, causing a periodic dimming of the star’s brightness, researchers kept an eye on these systems in the hopes of finding other transits Yang (2023). Identifying exoplanets became increasingly dependent on this methodology, particularly for those that were not initially found via the radial velocity method Bozza et al. (2016). Dedicated surveys were rapidly launched to perform "blind hunts," initially from space then from ground-based observatories Perryman (2018). The goal of these surveys was to find new planets only by using the periodic transit fingerprints observed in star light curves. With this method, scientists were able to find a wide variety of exoplanets without any prior awareness of their presence Chen (2023).
Figure 4: Exoplanets detection methods. fig 4.1 represents Number of exoplanets detected by Radial velocity method in previous years, fig 4.2 represents Number of exoplanets detected by Transit method in previous years, fig 4.3 represents Number of exoplanets detected by Gravitational Microlensing method in previous years, fig 4.4 represents Number of exoplanets detected by Direct Imaging method in previous years, fig 4.5 represents Number of exoplanets detected by Timing method in previous years, fig 4.6 represents Number of exoplanets detected by Astrometry method in previous years .
In the field of exoplanet research, transiting planets are especially significant Afanasev (2018). The light curves acquired during transits yield important details on planet sizes since the radii of the planets may be estimated due to the star’s periodic dimming of light Perryman (2018) Fischer et al. (2015). Through techniques like radial velocity measurements, astronomers can establish the transiting planet’s mass as well as its density, providing preliminary information about the planet’s composition Bozza et al. (2016) Chen (2023).
Figure 5: Representing the periodogram of exoplanet KIC 6922244
Around the end of the year of 2023, exoplanet research had undergone a major transformation. There were now close to 4153 transiting planets known, with the Kepler space observatory detecting nearly 2778 of them as reported by NASA. Kepler began operations in 2009 and has been a key player in the area of transiting exoplanet research, providing an abundance of data that has made significant effects Panahi et al. (2022) Chen (2023). A wide range of exoplanetary systems have been found as a result of the joint efforts of ground-based surveys Perryman (2018), such HAT and WASP, which revealed bright and easily detectable transiting systems, and Kepler’s numerous discoveries, particularly those of close-orbiting multi-planet systems Afanasev (2018). These systems contribute to our understanding of the diversity of planetary systems outside of our solar system by displaying a wide range of physical attributes and circumstances.
3.3 Direct imaging method
"Imaging" in the context of modern exoplanet research is the process of recognizing a planet as an individual source of light Perryman (2018) Chen (2023). This may happen as a result of the planet’s own thermal emission or light reflected from the host star. It is important that this imaging idea is not the same as reaching the planetary surface’s spatial resolution Bozza et al. (2016). Additionally, it is distinct from techniques that detect fluctuations in the total brightness Manfredi et al. (2020) of the planet and star system as a result of things like secondary eclipses, reflected light, or changes in the stellar spectrum Birkby (2018) throughout a transit. In 2005, the first imaging detections were reported, namely the identification of huge, youthful, self-luminous planets in broad orbits Bozza et al. (2016) Perryman (2018). The NASA exoplanet archive listed 44 of these discoveries as at the end of 2017 Perryman (2018). Prominent instances comprise planets discovered in well researched debris disk systems such as $\beta$ Pic and fomalhaut, in along with multi-planet systems like HR 8799 (which includes four planets) and LkCa 15 (containing two planets) Perryman (2018). There are certain deep learning models are used to inhance the capabilities of detection such as deep PACO Flasseur et al. (2024)
Researchers have been able to determine the orbital velocity of planets in systems such as $\beta$ Pic, HR 8799, and fomalhaut thanks to multi-epoch research. Over the past ten years, developments in imaging instrument technology Daglayan et al. (2023) Hebrard et al. (2018) have led to an emergence of cutting-edge devices specifically intended for exoplanet imaging Beuzit et al. (2006). VLT-SPHERE, Gemini-GPI Galicher et al. (2014), and Subaru-HiCIAO are a few examples Fischer et al. (2015). It is important to remember that these devices can only directly picture massive planets that are young, warm, and self-luminous at this time Perryman (2018). It is expected that with the next 30–40-meter telescopes, direct detection of point source pictures resulting from reflected starlight would become possible Bozza et al. (2016) Perryman (2018). On the other hand, picturing planets similar to earth is still a distant objective Chen (2023). It is crucial to take into account imaging Koechlin et al. (2005) at radio and X-ray wavelengths while studying star-planet systems. These other wavelength ranges may provide light on many facets of the interactions that take place between stars and their planets Flasseur et al. (2023). There is talk about the potential to expand imaging capabilities to X-ray and radio wavelengths, which implies a more thorough investigation of observational methods in exoplanet studies Rosa & do Nascimento Dias (2023) Daglayan et al. (2023).
3.4 Gravitational Microlensing
The general relativity-based observation method known as gravitational microlensing Perryman (2018). Using gravitational microlensing, scientists found 210 exoplanets by the end of 2023. This technique uses the gravitational potential of a foreground object (the lens) to bend light beams from a distant background object (the source) Bozza et al. (2016) Perryman (2018). Depending on how the source, lens, and observer are aligned, the outcome is a distorted image of the source, sometimes in many images Chen (2023) Beaulieu (2023). The effects of gravitational lensing can be observed, leading to the classification into various domains Jones (2008). Robust lensing, observable at the level of individual objects, is further classified into macro- and microlensing. Whereas microlensing deals with discrete, unresolved multiple visuals, macrolensing generates multiple resolved images, or "arcs." Weak lensing is a different regime that is statistically noticed without a clear resolution of individual objects. Specific circumstances within gravitational lensing are described by the introduction of specialized words. Mesolensing is the term used to characterize the astrometric displacement that accompanied a high-proper motion foreground star, whereas nanolensing is the lensing caused by planetary-mass objects Gaudi (2012). Relative motion between the source, lens, and observer causes time-varying amplification of pictures .
Depending on the characteristics of the radiation source and the lens, this amplification takes place across an assortment of intervals, from hours to years. A more complex light curve may be seen from the background source if the foreground lensing object is gravitationally complex, such a star with planets or a cluster of galaxies Wright & Gaudi (2012). This complexity results from the changing alignment geometry’s time-varying magnification. Significant discoveries have been made using gravitational microlensing Tamanini & Danielski (2019) in exoplanet exploration. In 2004, a 4MJ planet was first clearly detected at approximately 4 astronomical units (au) Perryman (2018). A 5M $\oplus$ planet was then found in 2006. An observation of a two-planet system in 2008 made it possible to quantify orbital motion during the lensing event. In 2015, the lens mass and microlens parallax were measured. In early 2018, evidence for planets flying freely was proposed. These results validate gravitational microlensing as an independent and potent probe of exoplanets, especially over an important mass and orbital radius region Bozza et al. (2016).
3.5 Astrometry
Table 3: Comparative analysis of methods
Detection method Principle Measurements Advantages speciality Drawbacks Equipement
Radial velocity Gravitational Wobble $m_{p},T,a$ large amount of planets can be observe at once, precision large planets, small cool stars its for only planets closer to observer Spectrometer
Transit photometry Light Dip $r_{p},T,a$ High senstivity, Large sample precise evaluation of properties, short period planets False positives,intersecting and sort period planets Photometer, TESS, Kepler, CoRoT
Direct Imaging Visual Observation $m_{p},T,a$ characterization , atmospheres large size and a, hot, planet close to observer needed stability, glare cornograph and infrared telescope
Gravitational Microlensing Gravitational Lensing $m_{p}$ Distant planets, galactic scale background towards center of galaxy, large a doubtful alignment, single trial Warsaw telescope, OGLE
Astrometry Stellar Wobble RA, Dec,.. High precision, Long baselines precise extremely hard Gaia, keck
The exact measurement of the motions and positions of different celestial objects, from bodies of the solar system and stars within our galaxy to, theoretically, entire galaxies and clusters, is known as astrophysicist. Determining stellar parallaxes and proper movements has historically been a major goal of astrometry, as these are essential for comprehending the characteristics of host stars Perryman (2018). In the present development, astrometry plays a new role: it seeks to identify the crosswise element of the displacement of a host star due to an orbiting planet’s gravity. Measures of radial velocity, which are sensitive to the concomitant movement of the photocenter along the line-of-sight, are closely associated with this dynamic manifestation in the sky Saffe et al. (2005).
It has been difficult to achieve high-accuracy astrometry; hipparcos and the fine guidance sensors on the hubble space telescope, for example, achieved accuracies of about 1 milliarcsecond (mas). Still, this degree of accuracy has only just touched the tip of the iceberg in terms of how much star position displacements caused by planets in orbit can be identified. By the end of 2023, astrometry had only found three exoplanets, two of which were 5MJ and 2.26MJ in mass, and one of which was a substellar partner with a mass 28 times that of jupiter (28MJ). These discoveries were recorded in the NASA exoplanet archive. Since the gaia Ranalli et al. (2018) mission was launched in 2013, the field of astrophysicists has seen a dramatic change Panahi et al. (2022). Over a billion stars with magnitudes up to V = 20–21 are being actively measured by gaia, with accuracies of about 20–25 microarcseconds (µas) for stars with a magnitude of 15 Perryman (2018). Our knowledge of planetary systems is predicted to be completely transformed by this astounding increase in precision Hodgkin et al. (2021) Panahi et al. (2022). Several thousand planetary systems with correct planet masses and absolute orbits free of the sin i ambiguity (the uncertainty surrounding the inclination angle of the planetary orbit) should be found thanks to gaia’s large dataset Panahi et al. (2022) Perryman (2018). Moreover, gaia will provide insightful data on planetary systems’ co-planarity, illuminating the configuration and behaviour of several planets within a system wikipedia (2022).
4 Comparative analysis of detection methods
Count of detected exoplanets by specific method is the proof of its importance. Every method of extrasolar planet detection works on different principles and scientific concepts Perryman (2018) Chen (2023). While transit photometry and radial velocity methods are playing main role for the majority detection. after discovery there are various parameters and properties to know for detecting habitable planet, which is possible through other methods Meunier et al. (2012). hence every detection method have its own advantages and disadvantages Fischer et al. (2015). The transit photometry method coveres wide range of stars but its false positive rate is high Tembhare et al. (2023). transit method works for star-planet systems, which have orbits aligned in such a way that, they can be visible from earth. those exoplanets should move in between host star and earth. Radial velocity method is highly prolific and works for many planets but they must be close to the spectator McArthur et al. (2004). Timing method works for limited materiality, but provide accurate data of exoplanets residing far from earth. Direct imaging also works for few stars but straight forward detection of exoplanets. Large orbital radii planets can be detected by gravitational microlensing, but here measurements alignments can’t be repeated and questionable Hebrard et al. (2018). The majority of precise searches employ multiple methodologies to collect multifaceted information on a certain planet. In particular exoplanet HD 86266 c alternative name TOI-652 c detected by primary transit method in 2020 and gave radius parameter details. Table 3 shows the comparative analysis of detection methods with important parameters.
5 Space missions for exoplanets
There are several missions are working towards exoplanets and there discoveries. these missions are running by various space agencies like NASA, ESA, JAXA, SETI etc. some missions ended up working and some are currently doing good work. here is the details of some space missions for exoplanets. The landscape of exoplanet exploration has been significantly shaped by a series of pioneering space missions, each contributing unique insights into the vast diversity of planetary systems beyond our solar neighborhood Galicher et al. (2014). Among these missions, TESS (Transiting Exoplanet Survey Satellite) Kane et al. (2020) and GAIA Ranalli et al. (2018) stand as ongoing endeavors,
Table 4: Space missions for exoplanet detection
Space missions for exoplanets
Mission Status
TESS Ongoing
GAIA Ongoing
Kepler completed (2013)
K2 Completed (2018)
CHEOPS On going
JWST On going
ARIEL Planned
RST Planned
TOLIMAN Planned
PLATO Planned
CoRot Completed(2006)
SPitzer Completed (2020)
tirelessly scanning the celestial expanse for transiting exoplanets and providing precise astrometric data for over a billion stars, respectively.There are various machine learning techniques Barboza et al. (2020) are used to study and detect exoplanets Jin et al. (2022). Table 4 shows the list of missions working on exoplanet wikipedia (2022) .
Kepler, although completed in 2013, left an enduring legacy by unveiling thousands of exoplanets through the meticulous scrutiny of stellar light curves. Its successor, K2, continued this mission, expanding our understanding of exoplanets until its completion in 2018. Ongoing missions such as CHEOPS (Characterizing Exoplanet Satellite) and JWST (James Webb Space Telescope) Alderson et al. (2023) hold promise for further revelations Agrawal et al. (2015). CHEOPS focuses on characterizing known exoplanets, scrutinizing their atmospheres and compositions, while JWST, a joint venture by NASA, ESA, and CSA, is poised to revolutionize infrared astronomy and deepen our understanding of exoplanet atmospheres Hebrard et al. (2018).
Looking ahead, a new wave of planned missions, including ARIEL (Atmospheric Remote-sensing Infrared Exoplanet Large-survey), RST (Roman Space Telescope), TOLIMAN, PLATO (PLAnetary Transits and Oscillations of stars), and more, signal a future of expanded horizons and enhanced capabilities in exoplanet detection and characterization. Reflecting on the completed missions, CoRot and Spitzer played pivotal roles, with CoRot contributing to the discovery of rocky exoplanets and Spitzer providing invaluable infrared observations until its completion in 2020. Together, these missions have propelled exoplanet science into a dynamic era, where each undertaking adds layers to the evolving narrative of the cosmos beyond our solar system wikipedia (2022).
6 Conclusion and future scope
A wide range of creative techniques used by astronomers and researchers have allowed the science of exoplanet detection to advance and diversify quite significantly. this detailed analysis of various exoplanet detection methods has illuminated the diverse approaches and technological advancements that have shaped our understanding of distant planetary systems. Each method, from the radial velocity and transit techniques to direct imaging, gravitational microlensing, astrometry, and pulsar timing, contributes uniquely to the growing catalog of known exoplanets. The precision of radial velocity measurements, the sensitivity of transits, and the potential for direct characterization through imaging collectively showcase the robustness of our exoplanet detection capabilities. Astrometry, with its emphasis on high precision and long baselines, offers complementary insights, especially for massive planets in close orbits.
Future work in exoplanet detection should strive towards enhancing the synergy between these methods. Integrating data from multiple detection techniques will not only strengthen the validity of individual findings but also provide a more comprehensive view of exoplanetary systems. Additionally, advancements in technology, particularly in the development of next-generation space telescopes and ground-based observatories, hold the promise of further refining existing methods and unveiling new avenues for exoplanet discovery. Continued collaboration between astronomers, astrophysicists, and instrument engineers will be crucial for pushing the boundaries of exoplanet science, ultimately enriching our understanding of the vast diversity of planetary systems beyond our solar neighborhood.